OpenStudy (anonymous):

solve the equation => In(e^18x)=9

OpenStudy (anonymous):

just remember that ln(e^a) = a*ln(e)

OpenStudy (anonymous):

what is the a tho?

OpenStudy (anonymous):

a parameter which can be everything, in our case a=18x

OpenStudy (anonymous):

so the equation is 18ln(e)^9 ??

OpenStudy (anonymous):

more like 18x * ln(e) = 9

OpenStudy (anonymous):

how much is ln(e) ?

OpenStudy (anonymous):

ln(e) = 1 that means our equation is 18x = 9 => x=1/2